Determine if the following conjecture is true. If not, give a counterexample.
The cubic power of an odd integer is odd.
Options:
true
false; 73 = 342
false; 93 = 728
OR
false; 53 = 126

Found 2 solutions by richard1234, Alan3354:
Answer by richard1234(7193)

Statement is true for all odd integers.

7^3 = 343, 9^3 = 729, 5^3 = 125 so the rest of those choices are all bogus.

Answer by Alan3354(69443)

Determine if the following conjecture is true. If not, give a counterexample.
The cubic power of an odd integer is odd.
Options:
true
false; 73 = 342
73 <> 342
false; 93 = 728
93 <> 728
OR
false; 53 = 126
53 <> 126
--------------
Do you mean to enter exponents?
Use ^ (Shift 6) for exponents.
===============
7^3 still not equal to 342, it's 343
=================
An odd integer to any integral power is and odd number.
(2n+1)^x = an odd number (n & x are integers)